how much water should be added to 30 liters of a 40 percent acid solution to reduce it to a 30 percent solution?
Answer by stanbon(75887)

how much water should be added to 30 liters of a 40 percent acid solution to reduce it to a 30 percent solution?
--------
Equation:
water + water = water
0.60*30 + x = 0.70(30+x)
-------
60*30 + 100x = 70*30 + 70x
30x = 10*30
x = 10 liters (amt. of water to add)
======================
Cheers,
Stan H.
======================